OpenStudy (anonymous):

A group of tourists went downstream a river and are supposed to return in 3 hours. How far can the tourists travel down the river if the speed of the current is 2mph, and the speed of the boat in still water is 18 mph?

OpenStudy (anonymous):

@Directrix @ash2326

OpenStudy (anonymous):

Ask @Hero !!

OpenStudy (anonymous):

@Agent_Sniffles @Luis_Rivera

OpenStudy (anonymous):

It depends if there is a waterfall or not.

OpenStudy (kropot72):

Let the distance = d Downstream speed = (18 + 2) mph Upstream speed = (18 - 2) mph \[time = \frac{distance}{speed}\] \[\frac{d}{18+2}+\frac{d}{18-2}=3\] Now solve for d to find the distance.